3x 13. Find the quotient of zx and 3 + x + 2. State any restrictions on the variable.
Answer #1

13. quotient of 9-22 and 23 + 2 +2 is, (33 = 19.) /19+pe+64) 3x (3+292 = (3-2) (344) (3+992 = 3x provided ex & -3 quotient isUsing simple polynomial division I solve the problem .